OpenStudy (anonymous):

A can do a piece of work in 20 days and B the same work piece of work in 40 days. if A and B work on alternate days, beginnin with A the work is finished during the course of the ........day when .......is working...... a.26th,A b.27th,A c.26th,B d.27th,B

OpenStudy (nowhereman):

OK, so the base data is: A is doing 1/20 per day and B is doing 1/40 per day. If they work daily alternating, in two days 1/20 + 1/40 = 3/40 is done in 2 days. Thus of two day units they need 40/3 = 13 1/3. Next take a closer look at the fourteenth unit. After 13 units, meaning 26 days 13*3/40 = 39/40 of the work is done and it's A's turn on the 27th day because A started. Also A will finish the work because it is 1/40 left and A can do 1/20, which is more. So answer b is correct.

OpenStudy (anonymous):

woow thats cool